a dam is built on a river to create a reservoir. the water level w in the reservoir is given by the equation w=3.5t+24 where t is the number of years the dam was constructed, and w is measured in feet.
graph the function

Answer by Theo(13342)

your function is w = 3.5 * t + 24.

to graph this function using graphing software, let y = w and x = t.

the function becomes y = 3.5 * x + 24.

y represents w which represents the water level, or the height of the water.

x represents t which represents the number of years since the dam was constructed.

you can graph it manually, but graphing software makes the job easier and it comes out neater as well as more accurate./

my software generated graph is shown below:

$$$

the values of y on the graph are generated from the formula of y = 3.5 * x + 24.

for example, the coordinate point (15,76.5) means the value of x is 15 and the value of y is 76.5

the formula would generate y = 3.5 * 15 + 24 which makes y = 76.5 when x is equal to 15.